OpenStudy (anonymous):

Can Anyone Help Me With This 16Question Test !?!?!

OpenStudy (anonymous):

Question 1 What are the slope and the y-intercept of this equation: y = -4x - 7 A. Slope = 4, y-intercept = 7 B. Slope = -4, y-intercept = -7 C. Slope = 4, y-intercept = -7 D. Slope = -4, y-intercept = 7 Question 2 What is the slope-intercept form of this equation? 4x + 2y = 12 A. y = -2x + 6 B. 2y = -2x + 12 C. 4x = -2y + 12 D. y = 2x + 6

OpenStudy (ashes.boom):

y = mx + b m = slope b = y intercept

OpenStudy (anonymous):

I think D

OpenStudy (lexi724):

@INeedSomeHelp! you have to post each question one at a time

OpenStudy (ashes.boom):

The 7 is not positive its negative okay

OpenStudy (anonymous):

I think 1 is {D} and I think 2 is {A}

OpenStudy (ashes.boom):

the first one is b

OpenStudy (ashes.boom):

second one is correct :) nice job

OpenStudy (anonymous):

Solve the equation for y: 10x + 5y = 30 A. 10x = -5y + 30 B. y = -10x + 30 C. y = -2x + 6 D. y = -2x + 30

OpenStudy (ashes.boom):

subtract the 10x from both sides

OpenStudy (ashes.boom):

then you have 5y = -10x + 30 and you can divide by 5 5y/5=y -10x/5= -2x 30/5=6 y = -2x + 6

OpenStudy (anonymous):

Thank youu , Which is an equation in point-slope form for the given point and slope? Point: (2, -5) Slope: -4/5 A. y - 5 = -4/5(x + 2) B. y + 5 = -4/5(x -2) C. y = -4/5 x - 2 D. y - 2 = -4/5(x - 5)

OpenStudy (ashes.boom):

okay so point slope form is y - y1= m(x - x1)

OpenStudy (ashes.boom):

y - (-5)= -4/5(x - 2)

OpenStudy (ashes.boom):

y + 5 = -4/5x and -4/5(2)

OpenStudy (ashes.boom):

y + 5 = -4/5x - 8/5

OpenStudy (anonymous):

So its {B} ? and What is the slope of the line that passes through these 2 points? (-2,3) and (2, -5) A. 2 B. -2 C. 1/2 D. -1/2

OpenStudy (ashes.boom):

Yea b sorry i kept going

OpenStudy (ashes.boom):

slope = y2-y1 / x2 - x1

OpenStudy (ashes.boom):

-5 - 3 / 2 - (-2) =?

OpenStudy (anonymous):

For some reason I think its a if not I think B

OpenStudy (ashes.boom):

-5 -3 is -8 and 2 - (-2) is 2 + 2 which is 4 so -8/4 is -2. its not a its b

OpenStudy (anonymous):

What does the slope of a line represent? A. Where the line crosses the y axis B. The ratio of the change in y values over the change in x values C. The ratio of the change in x values over the change in y values D. The height of the line

OpenStudy (ashes.boom):

what do you think it is

OpenStudy (anonymous):

C

OpenStudy (ashes.boom):

correct :) see you know it

OpenStudy (ashes.boom):

i will still be online but ihave to go to a class so i might not be able to answer ill try tho

OpenStudy (anonymous):

Okkk and thank you

OpenStudy (ashes.boom):

of course :) no problem